OpenStudy (anonymous):

Find the derivative of x-4 all under the square root sign. Using f(a+ delta x) -f(a) all over delta x.

OpenStudy (anonymous):

\[f(a)=\sqrt{a-4}\\ f'(a)=\lim_{\Delta x\to0}\frac{f(a+\Delta x)-f(a)}{\Delta x}\\ f'(a)=\lim_{\Delta x\to0}\frac{\sqrt{(a+\Delta x) -4}-\sqrt{a-4}}{\Delta x}\] Multiply the numerator and denominator by the conjugate of the numerator; that is, by \[\frac{\sqrt{(a+\Delta x) -4}+\sqrt{a-4}}{\sqrt{(a+\Delta x) -4}+\sqrt{a-4}}\] Then simplify.

OpenStudy (anonymous):

Does it come out to be 1 over 2 times the square root of x-4?

OpenStudy (anonymous):

Yes, that's right. I used an \(a\) instead of \(x\) above to follow the definition you're using.

OpenStudy (anonymous):

okay! Thank you so much!

OpenStudy (anonymous):

You're welcome.